If you have 42 coins totaling $7.20. and only have Dimes and Quarters, how many of each coin do you have?
Answer by jorel1380(3719)

d+q=42
10d+25q=720
10d+10q=420
15q=300
q=20
d=22
You have 22 dimes and 20 quarters..